# Deep learning modelling of manufacturing and build variations on multi-stage axial compressors aerodynamics

### v3

@@ 2.2 The C(NN)FD framework
The tip clearance and surface roughness values, as well as the relevant geometry design parameters of a specific build are fed as input to C(NN)FD, which predicts the 2D contours for all variables at the locations of interest. The outputs are mass-flow averaged to obtain the relevant radial profiles and 1D averages, which are then used to calculate stage-wise performance first, and then the overall performance using relevant thermodynamic equations. As the impact of tip clearance and surface roughness variations depends on the aerodynamic loading distribution of a given aerofoil, the methodology should be generalisable to different designs. Therefore, C(NN)FD requires the geometries associated with each clearance value as input. This is achieved by considering a series of design parameters to describe the geometry of each blade, such as stagger angle, camber angle, maximum thickness, etc. Since we consider only a single compressor design, the geometry parameters are fixed. In future work we will investigate the effect of geometrical variations and different designs.
2.2.1 Pre-processing
@@
The input of the network consists of two separate tensors. The first input is a tensor of size $(2\times 22)$, where the dimensions represent the input variables for each airfoil, respectively tip clearance and surface roughness, and the number of blade rows in the compressor. The second input is a tensor of size $(5\times 22\times 8)$ with the blade geometry design parameters. The dimensions represent the number of design variables, number of rows and number of radial sections. The compressor consists of 22 blade-rows, and aim of the network is to predict the flow field downstream of each of these blade-rows, at the corresponding 22 inter-row locations. However, for compressor aerodynamic analysis, often two additional locations are considered to calculate the overall performance: ”Compressor Inlet” and ”Compressor Outlet”, respectively upstream of the first row and downstream of the last row. This brings the number of axial locations for the predictions to 24. Therefore, the input tensor is manipulated by duplicating the input data for the first row at the beginning of the tensor, and the input data for the last row at the end of the tensor, bringing the first dimension of the tensor to 24. The input geometry tensor includes 8 radial sections at which the design parameters are specified, but this number can vary depending on the aerodynamic design philosophy for the compressor design of interest. Therefore, the radial sections are interpolated to computational grid, bringing input geometry shape to $(5\times 24\times 64)$. The geometry definition is then copied in the tangential direction to match the computational grid, defining a tensor of shape $(5\times 24\times 64\times 64)$. Likewise, the input tensor containing the tip clearance and surface roughness variables of shape $(2\times 24)$ is copied in the radial and tangential direction, to give a tensor of shape $(2\times 24\times 64\times 64)$. This allows to map the 2D and 3D input tensors to the 4D output tensor required for the flow field predictions. The two tensors are then added, resulting in a $(7\times 24\times 64\times 24)$ tensor, using a channel first convention. The number of channels is now representative of the input variables, including: tip clearance, surface roughness, inlet metal angle, outlet metal angle, maximum thickness, chord and pitch. The number of input variables could increase if considering more manufacturing and build variations. Therefore, a convolution is adopted to change the number of channels from the number of input design variables, to the number of output flow variables, which in this instance is 6. This results in a tensor of shape $(6\times 24\times 64\times 64)$ matching the requirements for the output flow field predictions. If more variables were required for the flow field predictions, it would only be required to extract the relevant data from the CFD solution and to increase the number of channels accordingly.
@@ 2.2.2 Neural Network architecture
The architecture shown in Figure 3, is a variant of the one originally proposed by the authors in previous work Bruni2023 . Following a pre-processing step for the input variables, the model is based on a 3D U-Net architecture with double residual convolutional blocks. Each convolutional block consists of a 3D convolution layer, followed by batch normalization and an ELU activation function. A residual connection is implement from after the first convolution to before the second activation function. The use of residual connections improves convergence and makes this type of architecture scalable to larger models required for full compressor applications. The number of axial locations is fixed to 24, and each down-sampling convolutions and up-sampling transposed convolutions uses a stride of $(1,2,2)$, resulting in a bottleneck section with tensors of shape $(384\times 24\times 1\times 1)$ after six layers. Both up and down samplings are followed by a double convolutional block at each layer. The output of the network is a tensor of size $(6\times 24\times 64\times 64)$, which encompasses the entire flow field. The encoding section followed by the decoding section, along with the skip connections, enables the network to predict both low-level and high-level features in the flow field.
@@ 2.2.3 Training
The network is trained using the AdamW optimizer, with Huber loss function. The dataset was divided into three sets: 70% training data, 20% validation data, and 10% hold-out data. The validation set was utilized to evaluate model performance and perform hyper-parameter tuning, while the hold-out set was reserved only for a final assessment. This ensures that the results are generalisable to unseen cases, and will lead to comparable performance also when deployed in a production environment. The dataset is split in a stratified manner, dividing the data in 10 bins using a quantile-based discretization function on the target overall efficiency. This ensures that each dataset has a comparable distribution. It should be noted that the dataset already contains out-of-tolerance cases and that no further out-of-distribution data outside of those limits is expected for the engineering application of interest. For further robustness, the training is performed using 5 different values of the random seed used for the dataset split, batch selection and initialization. The results are then presented providing the mean and the standard deviation of the losses. For instance, the final model showed a training loss of $(1.29\pm 0.16)\cdot 10^{-3}$ and a test loss of $(1.96\pm 0.40)\cdot 10^{-3}$ (Huber Loss $\pm$ standard deviation). An example of the training curve for one seed, starting from a random initialization in shown in Figure 4. The training runs generally converged within 500 epochs, with a batch size of 20 and an initial learning rate of 0.01. A learning rate scheduler was implemented, which halved the learning rate with a patience of 20 epochs. To avoid overfitting, early stopping with a patience of 50 epochs was implemented. The training time for a new model from scratch is around 3 hours on a single Tesla K80 GPU. This is reduced to less than 1 hour when re-training the network when data becomes available, thus making the architecture scalable for industrial applications. The execution time for inference is less than 1 second, making the predictions of the deployed model effectively real-time for the application of interest.
@@ 2.2.4 Model Comparison
The novelty of the C(NN)FD framework has been discussed in Section 1.4, highlighting how the appropriate pre-processing of the CFD solution and the usage of the relevant physical equations allow for the usage of CNNs for flow field predictions of complex 3D turbomachinery flows. However, the complexity of the regression problem will increase with the number of input features considered, due for example to: number of stages in the compressor and associated geometry parametrization, manufacturing variations considered (tip clearance, surface roughness, geometry variations or more) and operating conditions. Therefore, the complexity of the machine learning problem distinguishes this work from similar ones in literature Bruni2023 . For comparison, three different machine learning models are considered. The first one, Double Convolution consists only two convolution layers, each of which with 3D convolution, batch normalization, and ELU activation function. The second model, Standard U-Net is a 3D implementation of the U-Net, comparable to the ones used in previous literature for 2D airfoil flow predictions Thuerey2020 . The C(NN)FD is the architecture presented in this paper, and discussed in more details in Section 2.2.2. The main difference between the last two is that C(NN)FD includes residual connections within the double convolution blocks. Moreover, Standard U-Net uses a constant stride of $(2,2,2)$ across the three dimensions when downsampling and upsampling, which due to the target size of $(6\times 24\times 64\times 64)$ limits the width of the U-Net to three layers. In C(NN)FD the number of axial locations is instead fixed to 24 by using a stride of $(1,2,2)$, which allows for a wider U-Net with a total of six layers. The test loss of Double Convolution is $10.3x$ higher than C(NN)FD, while Standard U-Net is $3.28x$ higher. The accuracy achieved by these two architectures would not be sufficient to be used for the engineering application of interest. The improved accuracy of C(NN)FD is due to the higher number of trainable parameters and computational cost for C(NN)FD. However, more complex models such as U-Net++ Zhou2018 did not lead to an improvement in the accuracy of the predictions. Regardless, they will be considered in future work for more complex datasets, such as the ones including also geometrical variations or different operating conditions.
@@
Standard U-Net Thuerey2020 $3.19\cdot 10^{5}$ $2.3$ $5.40\cdot 10^{-3}$ $5.67\cdot 10^{-3}$
@@ 3 Results
In this section, the predicted flow field generated by C(NN)FD is compared to the CFD ground truth for the worst case scenario, where the largest discrepancy between the two was found in the hold-out set. A comparison between the ground truth on the left and the predictions on the right provides an overview of the flow field differences. The results are presented for Axial Velocity $V_{x}$ in Figure 5, Figure 6 and Figure 7 at Stage-1, Stage-5 and Stage-10. The relative error between the ground truth and the predictions at each mesh node are shown in Figure 8, Figure 9 and Figure 10 for Axial Velocity $V_{x}$ at Stage-1, Stage-5, and Stage-10. All the other stages and variables exhibit similar level of agreement and are not presented for conciseness.
@@
All values presented are non-dimensionalized with respect to the results obtained from the baseline. The Span [%] y-axis represents the radial direction in the gas path, ranging from the hub surface at the bottom to the casing surface at the top. The $\theta[\%]$ x-axis represents the circumferential direction in the gas path, with a range between 0 and 1 representing one passage. Each row corresponds to a specific number of blades, and thus the gas path can be divided into a number of passages equal to the blade count. The predictions generated by C(NN)FD exhibit excellent agreement with the ground truth, with the primary flow features being accurately reproduced. Only minor differences are observed in regions with high gradients, but well within the numerical uncertainty of the CFD solver. Due to operational requirements, the aerodynamic design of axial compressor leads to greatly different flow features across the different stages. For instance, front stages tend to operate at transonic conditions, with strong shock-waves interacting with the tip leakage flows. Conversely, the rear stages are subsonic, with the flow field dominated by end-wall flow features. The proposed architecture is demonstrated to be able to predict the flow field within the compressor for both transonic and subsonic stages, even with significantly different flow features as seen for instance by comparing Figure 5 and Figure 7. The highest local error is in the order of magnitude of $0.1\%$, which is well within the numerical error of the CFD solver. The errors are generally localised to regions with steep gradients and to only few nodes in the mesh, being therefore effectively negligible from an engineering perspective. When performing mass-flow averaging to obtain the radial profiles, the discrepancies between the ground truth and predictions are even less significant. Figure 11, Figure 12 and Figure 13 indicate that the radial profiles for C(NN)FD and CFD are effectively overlapping. The results are presented for Stage-1, Stage-5 and Stage-10. All the other stages and variables exhibit similar level of agreement and are not presented for conciseness. It is notable how even the end-wall regions, characterized by steep gradients, are well-resolved by C(NN)FD. This confirms that the localized errors found in the 2D contours prediction do not significantly affect the results for the radial profiles.
Figure 5: $V_{x}$ contour comparison between CFD and C(NN)FD predictions: Stage 1 - Rotor outlet and Stator outlet
Figure 6: $V_{x}$ contour comparison between CFD and C(NN)FD predictions: Stage 5 - Rotor outlet and Stator outlet
Figure 7: $V_{x}$ contour comparison between CFD and C(NN)FD predictions: Stage 10 - Rotor outlet and Stator outlet
Figure 8: $V_{x}$ error comparison between CFD and C(NN)FD predictions: Stage 1 - Rotor outlet and Stator outlet
Figure 9: $V_{x}$ error comparison between CFD and C(NN)FD predictions: Stage 5 - Rotor outlet and Stator outlet
Figure 10: $V_{x}$ error comparison between CFD and C(NN)FD predictions: Stage 10 - Rotor outlet and Stator outlet
@@
Figure 11: Radial profiles comparison between CFD and C(NN)FD predictions: Stage 1 - Rotor outlet and Stator outlet
@@
Figure 12: Radial profiles comparison between CFD and C(NN)FD predictions: Stage 5 - Rotor outlet and Stator outlet
@@
Figure 13: Radial profiles comparison between CFD and C(NN)FD predictions: Stage 10 - Rotor outlet and Stator outlet
The radial profiles are then mass-flow averaged to obtain the relevant 1D averages. An overview of the 1D average at each inter-row location is shown for Total Pressure $P_{t}$ in 13(a), Total Temperature $T_{t}$ in 13(b), Axial Velocity $V_{x}$ in 13(c) and Density $\rho$ in 13(d). Only these variables are presented as required to calculate the stage-wise and overall performance. However, similar agreement was also found for the remaining variables.
@@
The 1D averages are then utilized to compute the stage-wise performance as shown in 14(a). Predicting the stage-wise performance in addition to the overall performance is critical for implementation of the model in the gas turbine build process, as it provides detailed information on which blade rows should be targeted by corrective measures. While the agreement is excellent, minor discrepancies between predictions and ground truth are noticeable. C(NN)FD predicts accurately the trend for each stage, relative to the baseline, with the variations being slightly over-predicted or under-predicted depending on the stage. The large drop in Stage-10 efficiency could also potentially affect the compressor exit diffuser performance, which will also be considered in future work.
@@
The overall performance predictions are presented in Table 2, 15(a) and 15(b). An excellent agreement is observed for all overall performance parameters, with a coefficient of determination $R^{2}$ close to 1 for all variables. This indicates that the proposed model can accurately describe most of the variance present in the dataset. Moreover, the Mean-Absolute-Error for each variable is smaller than the known uncertainties of the CFD ground truth results and significantly smaller than the dataset’s range. If required, the predictions could potentially be further improved by providing training data including more extreme conditions, where it was deemed relevant from an industrial perspective.
@@
The C(NN)FD predictions remain accurate even for the case with the lowest efficiency in the hold-out set, demonstrating the model’s capability to handle challenging aerodynamic conditions typical of large clearances and surface roughness. The largest discrepancy shows a difference of less than $0.05$ percentage points in overall efficiency. This discrepancy is negligible for overall performance predictions, and demonstrates how the proposed architecture reliably predicts even the most challenging cases. This is a particularly challenging case aerodynamically, as high and low efficiency and pressure ratio stages are alternated, as shown in 14(b) . This would result for example in strong tip leakage flows leading to high losses and low efficiency in the blade rows with large clearances. Those would affect the downstream blade rows with tight clearance, whose potentially high efficiency will be affected by the unfavorable upstream conditions, leading to complex inter-row interactions and stage re-matching. This demonstrates how C(NN)FD can be used as a predictive tool to assess the performance of a given engine build, without the need for computationally expensive CFD analysis, and potentially avoiding re-testing for engines which would not meet the contractual performance requirements. Moreover, this tool can be used for a selective build process, where different parts available in stock can be selected based on the expected overall performance. The focus of this work was on compressor performance at design points conditions, while future work will extend the assessment to more off-design conditions and surge margin predictions.

### v4

@@ 2.3 The C(NN)FD framework
The tip clearance and surface roughness values, as well as the relevant geometry design parameters of a specific build are fed as input to C(NN)FD, which predicts the 3D flowfield, consisting of the 2D contours for all variables at the axial locations of interest. The outputs are mass-flow averaged to obtain the relevant radial profiles and 1D averages, which are then used to calculate stage-wise performance first, and then the overall performance using relevant thermodynamic equations. As the impact of tip clearance and surface roughness variations depends on the aerodynamic loading distribution of a given aerofoil, the methodology should be generalisable to different designs. Therefore, C(NN)FD requires the geometries associated with each clearance value as input. This is achieved by considering a series of design parameters to describe the geometry of each blade, such as stagger angle, camber angle, maximum thickness, etc. Since we consider only a single compressor design, the geometry parameters are fixed. In future work we will investigate the effect of geometrical variations and different designs.
2.3.1 Feature Engineering
@@
The input of the network consists of two separate tensors. The first input is a tensor of size $(2\times 22)$, where the dimensions represent the input variables for each airfoil, respectively tip clearance and surface roughness, and the number of blade rows in the compressor. The second input is a tensor of size $(5\times 22\times 8)$ with the blade geometry design parameters. The dimensions represent the number of design variables, number of rows and number of radial sections. The compressor consists of 22 blade-rows, and the aim of the network is to predict the flow field downstream of each of these blade-rows, at the corresponding 22 inter-row locations. However, for compressor aerodynamic analysis, often two additional locations are considered to calculate the overall performance: ”Compressor Inlet” and ”Compressor Outlet”, respectively upstream of the first row and downstream of the last row. This brings the number of axial locations for the predictions to 24. Therefore, the input tensor is manipulated by duplicating the input data for the first row at the beginning of the tensor, and the input data for the last row at the end of the tensor, bringing the first dimension of the tensor to 24. The input geometry tensor includes 8 radial sections at which the design parameters are specified, but this number can vary depending on the aerodynamic design philosophy for the compressor design of interest. Therefore, the radial sections are interpolated to computational grid, bringing input geometry shape to $(5\times 24\times 64)$. The geometry definition is then copied in the tangential direction to match the computational grid, defining a tensor of shape $(5\times 24\times 64\times 64)$. Likewise, the input tensor containing the tip clearance and surface roughness variables of shape $(2\times 24)$ is copied in the radial and tangential direction, to give a tensor of shape $(2\times 24\times 64\times 64)$. This allows to map the 2D and 3D input tensors to the 4D output tensor required for the flow field predictions. The two tensors are then added, resulting in a $(7\times 24\times 64\times 24)$ tensor, using a channel first convention. The number of channels is now representative of the input variables, including: tip clearance, surface roughness, inlet metal angle, outlet metal angle, maximum thickness, chord and pitch. The number of input variables could increase if considering more manufacturing and build variations. Therefore, a convolution is adopted to change the number of channels from the number of input design variables, to the number of output flow variables, which in this instance is 6. This results in a tensor of shape $(6\times 24\times 64\times 64)$ matching the requirements for the output flow field predictions. If more variables were required for the flow field predictions, it would only be required to extract the relevant data from the CFD solution and to increase the number of channels accordingly. One advantage of this architecture compared to others in literature, is that all the 6 variables of interest are predicted by a single model, instead of having to re-train the model separately for each variable of interest.
@@ 2.3.2 Neural Network architecture
The architecture shown in Figure 3, is a variant of the one originally proposed by the authors in previous work [2]. Following a pre-processing step for the input variables, the model is based on a 3D U-Net architecture with double residual convolutional blocks. Each convolutional block consists of a 3D convolution layer, followed by batch normalization and an ELU activation function. A residual connection is implement from after the first convolution to before the second activation function. The use of residual connections improves convergence and makes this type of architecture scalable to larger models required for full compressor applications. The number of axial locations is fixed to 24, and each down-sampling convolutions and up-sampling transposed convolutions uses a stride of $(1,2,2)$, resulting in a bottleneck section with tensors of shape $(384\times 24\times 1\times 1)$ after six layers. Both up and down samplings are followed by a double convolutional block at each layer. The output of the network is a tensor of size $(6\times 24\times 64\times 64)$, which encompasses the entire flow field. The encoding section followed by the decoding section, along with the skip connections, enables the network to predict both low-level and high-level features in the flow field.
@@ 2.3.3 Training
The network is trained using the AdamW optimizer, with Huber loss function. The dataset was divided into three sets: 70% training data, 20% validation data, and 10% hold-out data. The validation set was utilized to evaluate model performance and perform hyper-parameter tuning, while the hold-out set was reserved only for a final assessment. This ensures that the results are generalisable to unseen cases, and will lead to comparable performance also when deployed in a production environment. The dataset is split in a stratified manner, dividing the data in 10 bins using a quantile-based discretization function on the target overall efficiency. This ensures that each dataset has a comparable distribution. It should be noted that the dataset already contains out-of-tolerance cases and that no further out-of-distribution data outside of those limits is expected for the engineering application of interest. For further robustness, the training is performed using 5 different values of the random seed used for the dataset split, batch selection and initialization. The results are then presented providing the mean and the standard deviation of the losses. For instance, the final model showed a training loss of $(1.29\pm 0.16)\cdot 10^{-3}$ and a test loss of $(1.96\pm 0.40)\cdot 10^{-3}$ (Huber Loss $\pm$ standard deviation). An example of the training curve for one seed, starting from a random initialization in shown in Figure 4. The training runs generally converged within 500 epochs, with a batch size of 20 and an initial learning rate of 0.01. A learning rate scheduler was implemented, which halved the learning rate with a patience of 20 epochs. To avoid overfitting, early stopping with a patience of 50 epochs was implemented. The training time for a new model from scratch is around 1 hour on a single Tesla T4 GPU. This is reduced to less than 20 minutes when re-training the network when data becomes available, thus making the architecture scalable for industrial applications. The execution time for inference is less than 1 second, making the predictions of the deployed model effectively real-time for the application of interest.
@@ 2.3.4 Model Comparison
The novelty of the C(NN)FD framework has been discussed in Section 1.3, highlighting how the appropriate pre-processing of the CFD solution and the usage of the relevant physical equations allow for the usage of convolutional neural networks for flow field predictions of complex 3D turbomachinery flows. Regardless, the complexity of the regression problem will increase with the number of input features considered, due for example to: number of stages in the compressor and associated geometry parametrization, manufacturing variations considered (tip clearance, surface roughness, geometry variations or more) and operating conditions. Therefore, the complexity and dimensionality of the regression problem distinguishes this work from previous literature [2]. For comparison, three different machine learning models are considered. The first one, Double Convolution consists only two convolution layers, each of which with 3D convolution, batch normalization, and ELU activation function. This is considered just as a reference for what the simplest architecture that would be appropriate for the problem of interest. The second model, U-Net is a 3D implementation of the standard U-Net, comparable to the ones used in previous literature for 2D airfoil flow predictions [13], and re-implemented for a 3D formulation. The C(NN)FD is the architecture presented in this paper, and discussed in more details in Section 2.3.2. The main difference between the last two is that C(NN)FD includes residual connections within the double convolution blocks. The use of residual connections [25] was found to significantly improve the training for deeper networks, as it facilitates the propagation of the information across the network. Moreover, standard 3D U-Nets use a constant stride of $(2,2,2)$ across the three dimensions when downsampling and upsampling, which due to the flow-field target dimension of $(6\times 24\times 64\times 64)$ limits the width of the U-Net to three layers. In C(NN)FD the number of axial locations is instead fixed to 24 by using a stride of $(1,2,2)$, which allows for a wider U-Net with a total of six layers. In addition to the more complex architecture, fixing the first dimension to $24$ as the number of axial locations, is physically more meaningful. The network can learn the complex relationships between the different inter-row locations, as the flow field at a given location is defined by the upstream and downstream conditions, as well as the input features representative of that blade-row (i.e. geometry parameters, tip clearance and surface roughness). The test loss of Double Convolution is $10.3x$ higher than C(NN)FD, while for the standard U-Net is $3.28x$ higher. The accuracy achieved by these two architectures would not be sufficient to be used for the engineering application of interest. The improved accuracy of C(NN)FD is due to the residual connections, more physically meaningful convolutions, and higher number of trainable parameters for C(NN)FD. However, more complex models such as U-Net++ [26] did not lead to a significant improvement in the accuracy of the predictions, and were not considered worth of the increase computational cost. Regardless, they will be considered in future work for more complex datasets, such as the ones including also geometrical variations or different operating conditions.
@@
U-Net $3.19\cdot 10^{5}$ $2.3$ $5.40\cdot 10^{-3}$ $5.67\cdot 10^{-3}$
@@ 3 Results
In this section, the predicted flow field generated by C(NN)FD is compared to the CFD ground truth for the worst case scenario, where the largest discrepancy between the two was found in the hold-out set. Rather than selecting a cherry-picked case, as often the case in literature, comparing the results for the worst performer allows to identify the shortcoming of the model with regards to their physical interpretation and understand which flow features are more challenging to model using supervised learning approaches.
3.1 Flow field predictions
@@
The results are presented for Axial Velocity $V_{x}$ in Figure 5, Figure 6, Figure 7, Figure 8 and Figure 9, Figure 10 at Stage-1, Stage-5 and Stage-10, with CFD predictions in the left, C(NN)FD predictions in the middle, and relative error between the ground truth and the predictions at each mesh node on the right. All the other stages and variables exhibit similar level of agreement and are not presented for conciseness. The effect of errors in the other variables (Total Pressure ($P_{t}$), Total Temperature ($T_{t}$), Tangential Velocity ($V_{t}$), Radial Velocity ($V_{r}$), and Density ($\rho$)) are embedded in the assesment of the stage-wise distributions and overall performance, as they are used to calculate them directly through mass-flow averaging and the relevant physical equations. The $V_{x}$ comparison presented is used to judge the quality of the machine learning model in predicting the relavant aerodynamic flow features. This includes for example the regions of low axial momentum associated with the airfoil wakes, the tip clearance flows and suction side separations. All values presented are non-dimensionalized with respect to the results obtained from the baseline. The Span [%] y-axis represents the radial direction in the gas path, ranging from the hub surface at the bottom to the casing surface at the top. The $\theta[\%]$ x-axis represents the circumferential direction in the gas path, with a range between 0 and 1 representing one passage. Each row corresponds to a specific number of blades, and thus the gas path can be divided into a number of passages equal to the blade count. The predictions generated by C(NN)FD exhibit excellent agreement with the ground truth, with the primary flow features being accurately reproduced. Only minor differences are observed in regions with high gradients, but well within the numerical uncertainty of the CFD solver. Due to operational requirements, the aerodynamic design of axial compressor leads to greatly different flow features across the different stages. For instance, front stages tend to operate at transonic conditions, with strong shock-waves interacting with the tip leakage flows. Conversely, the rear stages are subsonic, with the flow field dominated by end-wall flow features. The proposed architecture is demonstrated to be able to predict the flow field within the compressor for both transonic and subsonic stages, even with significantly different flow features as seen for instance by comparing Figure 5 and Figure 9. The highest local error is less than $0.05\%$, which is well within the numerical error of the CFD solver. The errors are generally localised to regions with steep gradients and to only few nodes in the mesh, being therefore effectively negligible from an engineering perspective.
Figure 5: $V_{x}$ contour comparison between CFD and C(NN)FD predictions: Rotor 1 outlet
Figure 6: $V_{x}$ contour comparison between CFD and C(NN)FD predictions: Stator 1 outlet
Figure 7: $V_{x}$ contour comparison between CFD and C(NN)FD predictions: Rotor 5 outlet
Figure 8: $V_{x}$ contour comparison between CFD and C(NN)FD predictions: Stator 5 outlet
Figure 9: $V_{x}$ contour comparison between CFD and C(NN)FD predictions: Rotor 10 outlet
Figure 10: $V_{x}$ contour comparison between CFD and C(NN)FD predictions: Stator 10 outlet
3.2 Derived variables predictions
While C(NN)FD is trained exclusively on the flow field results from CFD, the 2D contours predictions are also used as a stepping stone to calculate other variables that are more frequently used for the engineering assessments. Mass-flow averaging in the circumferential direction is used to obtain radial profiles of the variables, which are then mass-flow averaged to obtain a 1D average, which is used to calculate stage-wise performance and overall performance of the variables of interest, such as pressure-ratio ($PR$), and polytropic efficiency ($\eta_{p}$). When performing mass-flow averaging to obtain the radial profiles, the discrepancies discussed for the 2D contours between the ground truth and predictions are even less significant. Figure 11, Figure 12 and Figure 13 indicate that the radial profiles for C(NN)FD and CFD are effectively overlapping. The results are presented for Stage-1, Stage-5 and Stage-10. All the other stages and variables exhibit similar level of agreement and are not presented for conciseness. It is notable how even the end-wall regions, characterized by steep gradients, are well-resolved by C(NN)FD. This confirms that the localized errors found in the 2D contours prediction do not significantly affect the results for the radial profiles.
@@ 3.2 Derived variables predictions
Figure 11: Radial profiles comparison between CFD and C(NN)FD: Stage 1 - Rotor and Stator outlet
@@
Figure 12: Radial profiles comparison between CFD and C(NN)FD: Stage 5 - Rotor and Stator outlet
@@
Figure 13: Radial profiles comparison between CFD and C(NN)FD: Stage 10 - Rotor and Stator outlet
The radial profiles are then mass-flow averaged to obtain the relevant 1D averages. An overview of the 1D average at each inter-row location is shown for Total Pressure $P_{t}$ in 14(a), Total Temperature $T_{t}$ in 14(b), Axial Velocity $V_{x}$ in 14(c) and Density $\rho$ in 14(d). Only these variables are presented as required to calculate the stage-wise and overall performance. However, similar agreement was also found for the remaining variables. Generally the trends are well captured, with $P_{t}$ and $\rho$ showing only negligible differences compared to the CFD predictions. While errors in $T_{t}$ and $V_{x}$ are slightly more noticeable for the rear stages of the compressor, they are still within the numerical uncertainty of the CFD calculations. Interestingly, while small in absolute terms, the errors are found to progressively increase towards the rear stages of the compressor. This is physically justifiable, as the flow field at these locations will be heavily influenced by the upstream conditions. Therefore, negligible errors in the front stages will progressively propagate towards the rear stages, leading to less accurate predictions, due to the higher uncertainty associated with the upstream flow field conditions. Future work will focus on quantifying the uncertainty of the predictions at each inter-row location, establishing the relative source of uncertainty which could be due to either shortcomings of the modelling of the effect of the input features for a given blade row, or the effect of the upstream and downstream boundary conditions for a given axial plane.
@@
The 1D averages are then utilized to compute the stage-wise performance as shown in 15(a) for pressure ratio and 15(b) for polytropic efficiency. While the agreement is excellent, minor discrepancies between predictions and ground truth are noticeable. C(NN)FD predicts accurately the trend for each stage, relative to the baseline, with the variations being slightly over-predicted or under-predicted depending on the stage but being ultimately negligible.
@@
Finally, the overall performance predictions are presented in Table 2, 16(a) and 16(b). An excellent agreement is observed for all overall performance parameters, with a coefficient of determination $R^{2}$ close to 1 for all variables. This indicates that the proposed model can accurately describe most of the variance present in the dataset. Moreover, the Mean-Absolute-Error for each variable is smaller than the known uncertainties of the CFD ground truth results and significantly smaller than the dataset’s range. If required, the predictions could potentially be further improved by providing training data including more extreme conditions, where it was deemed relevant from an industrial perspective.
@@
The C(NN)FD predictions remain accurate even for the case with the lowest efficiency in the hold-out set, demonstrating the model’s capability to handle challenging aerodynamic conditions typical of large clearances and surface roughness. The largest discrepancy shows a difference of less than $0.05$ percentage points in overall efficiency. This discrepancy is negligible for overall performance predictions, and demonstrates how the proposed architecture reliably predicts even the most challenging cases. This is a particularly challenging case aerodynamically, as high and low efficiency and pressure ratio stages are alternated, as shown in 15(b) . This would result for example in strong tip leakage flows leading to high losses and low efficiency in the blade rows with large clearances. Those would affect the downstream blade rows with tight clearance, whose potentially high efficiency will be affected by the unfavorable upstream conditions, leading to complex inter-row interactions and stage re-matching. This demonstrates how C(NN)FD can be used as a predictive tool to assess the performance of a given engine build, without the need for computationally expensive CFD analysis, and potentially avoiding re-testing for engines which would not meet the contractual performance requirements. Moreover, this tool can be used for a selective build process, where different parts available in stock can be selected based on the expected overall performance. The focus of this work was on compressor performance at design points conditions, while future work will extend the assessment to more off-design conditions and surge margin predictions.
3.3 Framework explainability
Predicting the flow field at all the inter-row axial locations, as well as radial profiles, 1D averages and stage-wise performance in addition to the overall performance is critical for implementation of the model in the gas turbine build process. The key advantage is that the wealth of information available provides actionable data on which blade rows should be targeted by corrective measures. The proposed architecture has a significant competitive advantage compared to traditional surrogate models, which would directly predict the overall performance of the compressor, starting from a defined set of input variables. In addition to being less generalisable and requiring more data for training, traditional ”black-box” models have the disadvantage of lacking explainability. Once a given overall performance variation is identified, it’s possible to back trace the results to the stage-wise distributions, 1D averages, radial profiles and ultimately 2D contours. This would be of significant use for root cause analysis, in which a variation in performance due to a certain input features, can be addressed by identifying the aerodynamic drivers associated with it. Further applications would consist of surrogate models for design activities or optimisations, once the model is trained considering also geometrical variations.